OpenStudy (anonymous):

how do i find the domain for this function f(x)=-5x+4

OpenStudy (anonymous):

Its a straight line right? and a line continues left and right forever and domain is all the x values, does that help?

OpenStudy (anonymous):

but how do i find the domain for this function that's where i'm confused

OpenStudy (noelgreco):

Your problem is not understanding what the term "domain of a function" means. Google it as a first try.

OpenStudy (anonymous):

i did but it doesn't give me nothing

OpenStudy (phoenixfire):

The Domain of a function is what values on the x-axis does the function include. This of this function f(x)=1/x. the domain of this function is (-infinity, 0) (0, infinity) But NOT including 0; hence the open bracket ) (. if it was closed bracket [ or ] then it means INCLUDING. So for your function as a straight line, it includes all values of x from -infinity to infinity

OpenStudy (noelgreco):

It didn't give me nothing: http://www.khanacademy.org/math/algebra/algebra-functions/v/domain-of-a-function Watch and ask if you don't get it. Please include what you didn't get and what you did to try to understand it. You will learn nothing by just being given the answer.

OpenStudy (anonymous):

great so you got it! "didn't give me nothing" two negative make a positive! The answer is that the domain is all real numbers

OpenStudy (noelgreco):

Simply giving her the answer is destructive.

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

well that's what i was asking how to find it not for the answer :)

OpenStudy (anonymous):

okay so i watched the video but how do i define when it's real numbers and from when a number is not included?

OpenStudy (anonymous):

generally when its a number that is not included.. it is either a point that is not on the graph.. or it is a negative under a radical.. or a denominator of 0

OpenStudy (anonymous):

the domain is simply all of the x values that the equation can assume. so in this case.. -5x+4 is a line.. because it is in the form y = mx + b. therefore.. it can assume all values of x.

OpenStudy (anonymous):

oh okay i got it now i understand thank you Nameless

OpenStudy (anonymous):

no problem. a graphing calculator helps out greatly if you cannot draw the graph out for yourself.

OpenStudy (anonymous):

and it took me a minute to grasp the concept of range and domain :) Good luck

OpenStudy (noelgreco):

If the question you asked is typical of the material you're working on, you've still got material your instructor hasn't presented. Example: What is the domain of the following function? \[\frac{ x }{ (x-4)(x+4) }\] Since a +4 or -4 would make the denominator = 0 (division by zero is not defined) we have to eliminate those two numbers from the domain thus: \[D: (-\infty,-4) \cup(-4,4)\cup(4,\infty)\] It looks bad, but if you keep current with the lessons you'll get it.

OpenStudy (anonymous):

 this is what i was thinking would be the set notation am i kind of like in the right track or no